After a genetics experiment on 63 pea plants, the number of plants having certain 23 were tall:; characteristics was tallied, with the results given to the right. Answer parts (a) through (c). 27 had green peas; 44 had smooth peas; 8 were tall and had green peas; (a) Find the number of plants that were tall and had smooth peas. 18 had green peas and smooth peas 3 had all three characteristics; 6 had none of the characteristics. plants
Answer #1

Let the number of plants that were tall and had only smooth peas be x

Green peas which are not tall are = 27-8 = 19

Only green peas = 27-5-15-3= 4

Only smooth peas = 44-15-x-3=26-x

Only tall = 23-5-3-x = 15-x

Only green and smooth = 18-3 = 15

Only tall and green = 8-3 = 5

Only tall and smooth = x

6 had none characteristic, hence with characteristics = 63-6 = 57

Only tall + only green peas + only smooth peas + only green and smooth + only tall and green + only tall and smooth + all 3 = 57

15-x +26-x+4+15+5+x+3 = 57

68-x = 57

Hence x = 11

Therefore number of plants that were tall and had smooth peas

= 11+3 = 14
Answer #2

After a genetics experiment on

pea plants, the number of plants having certain characteristics was tallied, with the results given to the right. Answer parts (a) through

(c).